OpenStudy (jamesreed30):

Express the required measurement an exact form find the length of an arc of 40° in a circle with an 8 inch radius use 3.14 for pie and round to hundreds

OpenStudy (3mar):

Did not get it yet?

OpenStudy (jamesreed30):

No @3mar

OpenStudy (3mar):

so may I send the attach again?

OpenStudy (jamesreed30):

Yea

OpenStudy (3mar):

Look again and I am here to help. Will not be late for any help.

OpenStudy (jamesreed30):

Okay so I looked at the image

OpenStudy (3mar):

Can you get something?

OpenStudy (jamesreed30):

I'm confused,the question says a circle though.

OpenStudy (3mar):

an arc of 40° in a circle with an 8 inch radius means part of the circumference of the circle!

OpenStudy (jamesreed30):

Okay so I've never worked with a problem like this before so could you take me through the steps please?

OpenStudy (3mar):

Of course With my pleasure!

OpenStudy (3mar):

To convert from degree to radian: \[\theta_{radian}=\frac{ \pi }{ 180 }*\theta_{degree}\] then it would be \[\theta_{radian}=\frac{ \pi }{ 180 }*40=0.7\] ok?

OpenStudy (jamesreed30):

Ok so 0.7 is the answer?

OpenStudy (3mar):

No, man. It is just the angle (in radian) which would be plugged in the equation or the arc length got it?

OpenStudy (jamesreed30):

Oh ok

OpenStudy (jamesreed30):

So what's next

OpenStudy (3mar):

Plug it in the eq. in the pic I sent! What you got?

OpenStudy (3mar):

it is easy now I think

OpenStudy (jamesreed30):

I would do 8 x 0.7?

OpenStudy (3mar):

Correct! This yields to ??

OpenStudy (jamesreed30):

5.6?

OpenStudy (3mar):

Correct!

OpenStudy (3mar):

Any more questions?

OpenStudy (jamesreed30):

Yes actually! I am doing online homeschool so without s teacher this stuff can get confusing so I have some more questions

OpenStudy (3mar):

Any Help... Anytime.

OpenStudy (jamesreed30):

I just posted another one

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (jamesreed30):

No problem

OpenStudy (3mar):

Can we start in new post, please?